OpenStudy (anonymous):

If secx > 0 and tanx < 0 then: A. 270

OpenStudy (anonymous):

\(\sec x = \dfrac{1}{\cos x}\), so secant will be positive when cosine is positive and will be negative when cosine is negative. Which quadrants will we find positive or negative secant? Use similar thinking to figure out in which quadrants \(\tan x = \dfrac{\sin x}{\cos x}\) is positive/negative, and use that information to figure out the problem (only one of the four quadrants will satisfy the condition).

OpenStudy (anonymous):

B

OpenStudy (anonymous):

No I mean D

OpenStudy (anonymous):

Good job! That is correct.